OpenStudy (lgbasallote):

\(\Huge \color{maroon}{\mathtt{LGBARIDDLE!!!}}\) Do you know your additon? Ride the lgbariddle expedition! Simple arithmetic: 1 + 1 = 3 2 + 2 = 3 3 + 3 = 3 4 + 4 = 2 5 + 5 = 3 6 + 6 = 2 7 + 7 = 2 8 + 8 = 2 9 + 9 = 4 what is the solution? credit to @asnaseer >:))

OpenStudy (anonymous):

1?

OpenStudy (lgbasallote):

\(\Huge \color{red}{\mathtt{X}}\) @Ishaan94

OpenStudy (anonymous):

\[\Huge\mathrm{\color{orangered}{4}}\mathtt{\color{turquoise}{2}}\]

OpenStudy (lgbasallote):

\(\Huge \color{pink}{\mathbf{still}}\) \(\Huge \color{skyblue}{\mathtt{no}}\)

OpenStudy (anonymous):

any number all we define before is null

OpenStudy (anonymous):

5+5 = NULL

OpenStudy (lgbasallote):

\(\LARGE \color{gold}{\mathsf{ha}}\)\(\Huge \color{green}{\mathbb{ha}}\) \(\large \color{red}{\mathcal{ha}}\)

OpenStudy (anonymous):

1=2=3=4=5 =NULL

OpenStudy (lgbasallote):

no answers? /:) hehe

OpenStudy (anonymous):

:(

OpenStudy (zepp):

Got it, 5.

OpenStudy (lgbasallote):

\(\Huge \color{yellowgreen}{\mathtt{\ddot \smile}}\) @Ishaan94 @zepp no :P

OpenStudy (apoorvk):

If, right, then explain, if wrong, go down the drain. :D

OpenStudy (zepp):

Hesitating between 2 and 3 :/ Let's go for 3 :D

OpenStudy (lgbasallote):

lol if you know the solution you be no hesitate :P

OpenStudy (zepp):

Well, the solution is int(2.5) D:

OpenStudy (lgbasallote):

lol no

OpenStudy (zepp):

Pretty sure my thing works, lemme draw it

OpenStudy (moongazer):

10 ?

OpenStudy (anonymous):

4?

OpenStudy (lgbasallote):

no..no..and no..

OpenStudy (anonymous):

Just for a wild guess, 11?

OpenStudy (lgbasallote):

no :))

OpenStudy (anonymous):

Haha. That's an interesting problem, but my brain is way too tired to come up with anything :-)

OpenStudy (blockcolder):

5+5=x where x is the answer.

OpenStudy (lgbasallote):

hmmm good thinking...

OpenStudy (anonymous):

This better be a good one.

OpenStudy (anonymous):

I think it isss... 4.

OpenStudy (lgbasallote):

i hope so too fool :P asnaseer just taught it to me haha no @mattt9

OpenStudy (anonymous):

19?

OpenStudy (lgbasallote):

nope :D where are you getting these numbers anyway? lol

OpenStudy (anonymous):

1,2,3...∞

OpenStudy (anonymous):

l.h.s-r.h.s :D

OpenStudy (anonymous):

Okay, so it's not 1,2,3,4,5 right?

OpenStudy (lgbasallote):

i dont think tht's the formula//

OpenStudy (lgbasallote):

it's one of those

OpenStudy (anonymous):

It's 2!

OpenStudy (lgbasallote):

nope :P

OpenStudy (lgbasallote):

Hint: it's a riddle so real math is less likely :P

OpenStudy (anonymous):

But you have already turned down 1,3,4 and 5. I think it should be 3 though.

OpenStudy (lgbasallote):

lol i didnt turn down 3 =)))

OpenStudy (anonymous):

Aha I knew it! :D

OpenStudy (lgbasallote):

but what's the solution /:)

OpenStudy (anonymous):

is it 2?

OpenStudy (anonymous):

+1,-1,0-2,+2 :) <---- meh guess lols

OpenStudy (lgbasallote):

no and no

OpenStudy (anonymous):

6?

OpenStudy (lgbasallote):

no...it's 3...like fool said...im asking the solution now :P

OpenStudy (anonymous):

I'm thinking...dont rush me.

OpenStudy (lgbasallote):

maybe i should post more numbers hmm..

OpenStudy (apoorvk):

NOOOOOOOOOOOOOOOOOOOOOOO!!!!!!!!!!!!!!!!

OpenStudy (apoorvk):

okay, post more......... YESSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSS!!!!!!!!!!!!!!!!! :P

OpenStudy (anonymous):

The next term is 4 right?

OpenStudy (lgbasallote):

no.....10 + 10 = 2

OpenStudy (anonymous):

Okay this is a 42 problem.

OpenStudy (anonymous):

http://mathworld.wolfram.com/42.html

OpenStudy (anonymous):

5-2=3 (5) 7-4=3(7) 10-8=2 6-3=3 8-5=3 *like a boss*

OpenStudy (lgbasallote):

lol =))))) so that's what ishaan meant

OpenStudy (lgbasallote):

random numbers lol becca

OpenStudy (anonymous):

It's not...post your original question again..it makes perfect sense

OpenStudy (lgbasallote):

whered you get 5, 2, 7, 4 , etc

OpenStudy (anonymous):

can you just post the orignal...then I shall show you :D

OpenStudy (lgbasallote):

1 + 1 = 3 2 + 2 = 3 3 + 3 = 3 4 + 4 = 2 5 + 5 = ?

OpenStudy (lgbasallote):

show me then

OpenStudy (lgbasallote):

Hint 2: "simple" arithmetic haha

OpenStudy (anonymous):

okay 3 +2 =5 4+3=7 6+3=9 8+2=10 5+5=3 ( I saw the answer so I kinda cheated) then go back through the original and 5-2=3 7-4=3 10-8=2 6-3=3 8-5=3

OpenStudy (zepp):

becca, you kinda did the same thing but in the opposite way..?

OpenStudy (lgbasallote):

lol 6-3 and 8-5 =))) that's still not the formula

OpenStudy (anonymous):

darn, oh well hhahaa I have failed you

OpenStudy (lgbasallote):

should i tell the answer now...

OpenStudy (anonymous):

69 c;

OpenStudy (anonymous):

No.

OpenStudy (lgbasallote):

okay..

OpenStudy (anonymous):

Okay post it now! This better be a very good one :P

OpenStudy (lgbasallote):

that makes me nervous that it might not be good enough :P but what the heck it's a riddle...

OpenStudy (lgbasallote):

this was a two part riddle..the first is the "simple" arithmetic part...it means to use kindergarten arithmetic...that means 1 + 1 is not 2..but becomes 11 so... 1+1 = 11 2+ 2 = 22 3+3 = 33 4+4 = 44 5+5 = 55 6+6 = 66 7+7 = 77 8+8 = 88 9+9 =99 11 = ElEvEn <- 3 Es 22 = TwenTy Two <- 3 Ts 33 = ThirTy Three <- 3 Ts 44 = Forty Four <- 2 Fs 55 = FiFty Five <- 3 Fs 66 = Sixty Six <- 2 Ts 77 = Seventy Seven <- 2 S 88 = Eighty Eight <- 2 Es 99 = NiNety NiNe <- 4 Ns good enough @FoolForMath :troll:

OpenStudy (anonymous):

But I wouldn't have guessed it in a million years.

OpenStudy (anonymous):

There is not symmetry on the letters of repetition right?

OpenStudy (anonymous):

Okay got it! Good one :)

OpenStudy (anonymous):

Thanks for sharing!

OpenStudy (lgbasallote):

i dont know..it's just the number of repetitions of the first letter...asnaseer suggested it while he was solving a harder problem haha

OpenStudy (lgbasallote):

i hope i dont get banned for a riddle with a ridiculous solution @mattfeury *whistle*

OpenStudy (apoorvk):

Great one for outta the box thinking though, @lgbasallote and ofcourse @asnaseer !

Parth (parthkohli):

10 + 10 = 20